OpenStudy (anonymous):

sqrt x+3-2=2 solve for x

OpenStudy (wwwhhhaaattt?):

You cant leave me

OpenStudy (wwwhhhaaattt?):

Try

OpenStudy (wwwhhhaaattt?):

@jigglypuff314

OpenStudy (anonymous):

HELPP !! @OGangsta @Kate101 @LeeEtchison @LexiGirl777 @love10129151 @Pi123 @LadyM @phi @phi @highschoolmom2010 @Blu-Girl @rtracy @eighthourlunch @gibbs @Ivanloza96 @dan815 @Shadowgirl15 @Dmex @Rainbow_dashie @quickstudent @haleyelizabeth2017 @Frazfootballa @freckles @Godlovesme @bloopman

OpenStudy (wwwhhhaaattt?):

BLOOP NOT YET, DONT HELP

OpenStudy (anonymous):

whaa

OpenStudy (wwwhhhaaattt?):

@kdog55 TRY

OpenStudy (wwwhhhaaattt?):

Your not trying

OpenStudy (anonymous):

@LexiGirl777 can u help

OpenStudy (wwwhhhaaattt?):

NO ONE ANSWER

OpenStudy (wwwhhhaaattt?):

DONT HELP

OpenStudy (anonymous):

i am going to answer!

OpenStudy (anonymous):

It's logic. Add two to both sides. sqrt(x+3) -2 = 2 sqrt(x+3) = 4 Now, what's the only number that, when taken the square root of, gives you 4?

OpenStudy (wwwhhhaaattt?):

DONT ANSWER

OpenStudy (godlovesme):

lol @WWWHHHAAATTT?

OpenStudy (rainbow_dashie):

1+ x=2

OpenStudy (anonymous):

?????????I'm so confused on what's going on?????????

OpenStudy (wwwhhhaaattt?):

He just wants the answer

OpenStudy (anonymous):

ok and you don't want us to answer it???

OpenStudy (anonymous):

im a GIRL

OpenStudy (anonymous):

bye i am to confused.sorry @kdog55

OpenStudy (anonymous):

LOL

OpenStudy (godlovesme):

rofl

OpenStudy (anonymous):

@bloopman I got 2

OpenStudy (anonymous):

I'm trying to help you.

OpenStudy (anonymous):

ME 2 IM LOST BYE

OpenStudy (wwwhhhaaattt?):

Because ive been in 3 questions hes asked but ever time i tryed to tell him to try firsst he said he did but he wouldent show me

OpenStudy (anonymous):

its a girl

OpenStudy (rainbow_dashie):

Hush I'm helping him @kdog55 3-2= what

OpenStudy (wwwhhhaaattt?):

Rainbow Dont

OpenStudy (anonymous):

@Rainbow_dashie 1

OpenStudy (anonymous):

No, not when it is squared, but when it is taken the square root of.

OpenStudy (anonymous):

@bloopman i dont understand

OpenStudy (anonymous):

2^2 = 4, but sqrt(?) = 4

OpenStudy (wwwhhhaaattt?):

@kdog55 Try

OpenStudy (rainbow_dashie):

Okay so x+1+2

OpenStudy (wwwhhhaaattt?):

@Maria195

OpenStudy (anonymous):

@bloopman 16

OpenStudy (anonymous):

@WWWHHHAAATTT? shutup

OpenStudy (wwwhhhaaattt?):

User Disrespect is agianst CoC

OpenStudy (rainbow_dashie):

K bye I'm not even helping

OpenStudy (anonymous):

Correct, good job. Back to what we had earlier: sqrt(x + 3) = 4 Since 16 is the only thing that can betaken the square root of to get 4, that means x + 3 = 16. So, what is x in the equation x + 3 = 16?

OpenStudy (wwwhhhaaattt?):

Stop helping

OpenStudy (anonymous):

@bloopman is it 13?

OpenStudy (anonymous):

I'll help as I please babe

OpenStudy (anonymous):

@kdog55, yes. That's your answer x = 13

OpenStudy (anonymous):

thank you , for all your help, can u help wit another one @bloopman

OpenStudy (anonymous):

The strictly algebraic way to do this is like so: sqrt(x+3) - 2 = 2 sqrt(x+3) = 4 square both sides x+3 = 16 x = 13 Sure, Ill help with another

OpenStudy (anonymous):

\[10+\sqrt{v-4=6}\] @bloopman

OpenStudy (wwwhhhaaattt?):

If you need help with another Open a new question

OpenStudy (anonymous):

nobody is talking to you @WWWHHHAAATTT? , thank u for your NONSERVICE

OpenStudy (anonymous):

@WWWHHHAAATTT? that's not required.

OpenStudy (anonymous):

@kdog55, are you sure you typed that equation correctly? (besides the sqrt going over the equal sign, obviously) have you learned imaginary numbers?

OpenStudy (anonymous):

Yes im sure

OpenStudy (anonymous):

the sqrt sign is only over the v-4 @bloopman

OpenStudy (anonymous):

\[10+\sqrt{v-4}=6\] like this

OpenStudy (anonymous):

okay, so 10 +sqrt(v-4) = 6 first things first, subtract 10 from both sides sqrt(v-4) = -4 now, do u know what we can do from here?

OpenStudy (anonymous):

wait huh,

OpenStudy (anonymous):

10 - 6?

OpenStudy (anonymous):

no, 6 - 10

OpenStudy (anonymous):

we're isolating the square root

OpenStudy (anonymous):

ok -4

OpenStudy (anonymous):

right, so how do we solve v in sqrt(v-4) = -4

OpenStudy (anonymous):

add 4 both sides?

OpenStudy (anonymous):

we can't do that, because the 4 in the square root. we have to get it out of the square root (take your time, don't get frustrated.) i told you implicitly how to do this earlier

OpenStudy (anonymous):

square both sides ?

OpenStudy (anonymous):

yep. :D so what do u get for that?

OpenStudy (anonymous):

am i squaring -4? or what

OpenStudy (anonymous):

you're squaring both sides of the equation

OpenStudy (anonymous):

so you will square -4 on the right side

OpenStudy (anonymous):

idk how to do that, like what numbers

OpenStudy (anonymous):

is it v-4 = -16

OpenStudy (anonymous):

you just square sqrt(v-4) on the left = -4 so (sqrt(v-4))^2 = (-4)^2

OpenStudy (anonymous):

what happens when you square a square root?

OpenStudy (anonymous):

idk i got that and just added -16+4

OpenStudy (anonymous):

think about it. thw a

OpenStudy (anonymous):

*the square root of 16 is 4. 4^2 = 16. so what does that tell you about (sqrt(x))^2 for any x?

OpenStudy (anonymous):

im totally lost,

OpenStudy (anonymous):

its a perfect square

OpenStudy (anonymous):

alright. so we took the square root of something, then we squared it. now we GOT that original something. 2^2 = 4, sqrt(4) = 2. therefore (sqrt(4))^2 = 4 3^2 = 9, sqrt(9) = 3. therefore sqrt(9))^2 = 9

OpenStudy (anonymous):

and so on. so that means the square of a square root is equal to the number inside the square root.

OpenStudy (anonymous):

then whats my answer 16 or 4

OpenStudy (anonymous):

in context, this means (sqrt(v-4))^2 = = v-4. get it?

OpenStudy (anonymous):

we took the square of the square root of v-4 and got v-4

OpenStudy (anonymous):

so v-4 = 16

OpenStudy (anonymous):

yep! now just solve for v

OpenStudy (anonymous):

16+4?

OpenStudy (anonymous):

mhm.

OpenStudy (anonymous):

20 not the answer

OpenStudy (anonymous):

what are your options?

OpenStudy (anonymous):

there are none

OpenStudy (anonymous):

could be a no solution

OpenStudy (anonymous):

its either the answer or no solution